Find two consecutive whole numbers such that five times the smaller number plus three times the greater number makes 59
(Hint: let the numbers be x and (x+1)

Answer by ikleyn(52781)

.
Find two consecutive whole numbers such that five times the smaller number plus three times the greater number makes 59
(Hint: let the numbers be x and (x+1)
~~~~~~~~~~~~~~~

Following to the HINT, you can write an equation as you read the problem

    5x + 3*(x+1) = 59.


Simplify and find x

    5x + 3x + 3 = 59

    5x + 3x     = 59 - 3

      8x        =   56

       x        =   56/8 = 7.


ANSWER.  The numbers are 7 and 8.


CHECK.  5*7 + 3*8 = 35 + 24 = 59.    ! correct !

Solved.